# A weak regularity lemma for polynomials

### v1

@@ 1.3 The proof
Towards the proof of the weak regularity lemma, we obtain results on linear spaces of polynomials and forms (homogeneous polynomials) that may be of independent interest. In particular, we analyze subspaces of high rank polynomials (where the rank is allowed to grow with the number of spanning polynomials, i.e., the dimension of the space). More particularly, we analyze high-rank polynomial “pencils”. By a polynomial pencil we mean a set of the form $V\setminus U$ where $V$ is a linear space of polynomials and $U$ is a strict subspace. For comparison, the classical regularity lemmas for polynomials work only with the “trivial” pencil $V\setminus\{0\}$; or more concretely, all nonzero linear combinations of the parts $\mathbf{X}$ are required to have high rank. Obtaining a high-rank pencil, rather than a high-rank trivial pencil, turns out to be achievable with far better bounds. We call this result a “rank-regularity lemma”. Using it, the rest of our proof has, very roughly, the following logical structure (see Section 2 for all definitions):
$\mathbf{X}\text{ $t$-rank-regular}$ $\Leftrightarrow\operatorname{rk}(X_{1}+\operatorname{Sp}\mathbf{X}^{\prime})>tr$
$\Rightarrow|\operatorname{bias}(X_{1}+\operatorname{Sp}\mathbf{X}^{\prime})|<\epsilon q^{-k}$
@@
The proof moreover guarantees the additional required conditions, e.g., that $X_{1}$ has maximal degree. To translate from high rank to low bias as above, our proof relies on recent results on the relationship between structure and randomness for polynomials (see e.g. [14, 24, 4, 25]). Incidentally, these results are proved by going through the corresponding multilinear forms—which requires the characteristic of the field to be larger than the degree of the parts—by relating analytic rank and partition rank (defined by Gowers and Wolf [8] and by Naslund [26], respectively).
@@ 1.4 Organization
In Section 2 we state and prove our weak regularity lemma. The proof has two components: in Section 2.1, we prove a “simplified” version, which finds a small decompositions where every linear combination outside of a strict subspace has high rank; in Section 2.2 we show that a similar condition in terms of bias (rather than rank) suffices to get weak regularity, using Fourier-analytic techniques and known structure-vs-randomness results; finally, in Section 2.3, we combine the aforementioned pieces to prove the weak regularity lemma. In Section 3 we obtain the applications of the weak regularity lemma outlined above.
@@ 1.5 Preliminaries
We will frequently use boldface letters to emphasize that an object is a tuple; so for example, we will use $\mathbf{P}$, instead of $P$, when dealing with a tuple of polynomials (or a polynomial map). We will use the convention that for a tuple $\mathbf{X}=(X_{1},\ldots,X_{k})$ and a set $A$, the notation $\mathbf{X}\subseteq A$ is a shorthand for $X_{i}\in A$ for all $i$. For $\mathbf{P}\subseteq\operatorname{Poly}$ we denote $\deg(\mathbf{P})=\max_{i}P_{i}$.
@@ 2 Weak Regularity
Definition 2.1.
[weak $\epsilon$-regularity] Let $\mathbb{F}=\mathbb{F}_{q}$. We say that $\mathbf{X}\in\operatorname{Form}^{k}(\mathbb{F})$ is weak $\epsilon$-regular if $\deg(X_{1})=\deg(\mathbf{X})$, and for every $y\in\mathbb{F}^{k}$,
@@
Note that the probabilistic condition in Definition 2.1 implies that $\Pr[\mathbf{X}=\mathbf{y}]=(\alpha\pm\epsilon)q^{-k}$ for any $\mathbf{y}\in\mathbb{F}^{k}$ with $\Pr[\mathbf{X}^{\prime}=\mathbf{y}^{\prime}]=\alpha q^{-(k-1)}$ (recall $\mathbf{X}^{\prime}\in\operatorname{Form}^{k-1}$). Intuitively, this means $\mathbf{X}$ is distributed like $\mathbf{X}^{\prime}$ up to an error of $\epsilon$.
@@ 2.1 Rank-regularity
An important definition in the study of polynomials is that of rank, sometimes also called strength [1] or Schmidt rank [28]. Recall that a polynomial is reducible if it can be factorized as the product of two non-constant polynomials (all polynomials over the same field).
@@
Every $\mathbf{P}\in\operatorname{Poly}^{m}_{d}$ has a weak $t$-rank-regular decomposition of size at most $((2t+1)dm)^{2^{d}}$.
@@ 2.1.1 Linear spaces of polynomials
$V^{\uparrow}=\{X\in V|\deg(X)=\deg(V)\}.$
We will need the following closure property.
@@ 2.1.2 Proof of the rank-regularity lemma
Let $\mathbf{X^{0}}\in\operatorname{Form}^{m^{\prime}}$ be obtained from $\mathbf{P}\in\operatorname{Poly}^{m}$ by replacing each component of $\mathbf{P}$ with its homogeneous parts of degree $1,2,\ldots,d$. Note that $\mathbf{P}\subseteq\mathbb{F}[\mathbf{X^{0}}]$ and $m^{\prime}\leq dm$.
@@
For each $i$ starting with $i=0$, we construct $\mathbf{X^{i+1}}$ from $\mathbf{X^{i}}$ as follows:
  • Consider a minimal subspace $W$ of $\operatorname{Sp}(\mathbf{X^{i}})$, with a basis of forms $\mathbf{Y}\subseteq\operatorname{Form}^{r^{\prime}}$, satisfying $\mathbf{P}\subseteq\mathbb{F}[W]$. Then $\mathbf{P}\subseteq\mathbb{F}[\mathbf{Y}]$ is a minimal decomposition, and moreover, $\deg(\mathbf{Y})\leq\deg(W)\leq\deg(\mathbf{X^{i}})$, and $r^{\prime}=\dim W\leq\dim\operatorname{Sp}(\mathbf{X^{i}})\leq r_{i}$.
  • Now, if $\mathbf{P}\in\mathbb{F}[\mathbf{Y}]$ is $t$-rank-regular then set $\mathbf{X^{i+1}}=\mathbf{Y}$ and $s=i+1$, finishing the construction. Otherwise, $\mathbf{Y}$ is not $t$-rank-regular. We will next apply Lemma 2.9 on the forms in $\mathbf{Y}$ of maximal degree, $\ell:=\deg(\mathbf{Y})$. More precisely, we apply Lemma 2.9 on the tuple $\mathbf{Y^{*}}\in\operatorname{Form}_{\ell}^{r^{\prime}}$ obtained from $\mathbf{Y}$ be replacing every component of degree smaller than $\ell$ by an arbitrary component of degree $\ell$; since $\operatorname{Sp}(\mathbf{Y^{*}})=\operatorname{Sp}(\mathbf{Y})^{\uparrow}$, and both tuples are of the same size, we have from Corollary 2.8 that $\mathbf{Y^{*}}$ is also not $t$-rank-regular, as needed. Merging the resulting tuple of forms with the forms in $\mathbf{Y}$ of degree smaller than $\ell$, we obtain $\mathbf{X^{i+1}}\in\operatorname{Form}^{r_{i+1}}$ with $\mathbf{Y}\subseteq\mathbb{F}[\mathbf{X^{i+1}}]$, $\deg(\mathbf{X^{i+1}})<\deg(\mathbf{X^{i}})$, and $r_{i+1}\leq 2tr_{i}^{2}+r_{i}\leq(2t+1)r_{i}^{2}$.
To prove that this construction gives the claim above, we proceed by induction on $i\geq 0$. The induction base $i=0$ follows trivially from the construction of $\mathbf{X^{0}}$, as $\mathbf{P}\subseteq\mathbb{F}[\mathbf{X^{0}}]$ with $\deg(\mathbf{X^{0}})\leq d$, and $r_{0}\leq dm=(2t)^{0}(dm)^{1}$. For the induction step, we have, by the construction of $\mathbf{X^{i+1}}$ and the induction hypothesis, that $\mathbf{P}\subseteq\mathbb{F}[\mathbf{Y}]\subseteq\mathbb{F}[\mathbf{X^{i+1}}]$, that $\deg(\mathbf{X^{i+1}})\leq\deg(\mathbf{X^{i}})-1\leq d-(i+1)$, and that $r_{i+1}\leq(2t+1)r_{i}^{2}\leq(2t+1)((2t+1)^{2^{i}-1}(dm)^{2^{i}})^{2}=(2t+1)^{2^{i+1}-1}(dm)^{2^{i+1}}$, as needed. This completes the proof. ∎
@@ 2.2 Bias and weak regularity
We note that $|\operatorname{bias}(P)|$ is closely related to the statistical distance between the distribution of $P$ and the uniform distribution on $\mathbb{F}$ (see e.g. [3], Claim 33).
@@
Theorem 2.10 ([25], Corollary 8.5121212See also Remark 8.3 for the bound on $c_{d}$.).
@@
We also remark that one can replace the use of Theorem 2.10 in our proof by earlier bounds of Milićević [24] or Janzer [14], and still obtain reasonable, albeit somewhat worse bounds in Theorem 2.2.
@@ 2.3 Putting everything together
Take $t$ such that the bound in Theorem 2.10 satisfies $q^{-c_{d}tk/\mathcal{L}_{\mathbb{F}_{q}}(tk)}\leq\epsilon q^{-k}$ for every $k$; note that $t=2^{d^{1+o(1)}}(r+1)^{1+o(1)}$ suffices. [^15] In fact, this choice of $t$ suffices even if the right hand side is $\epsilon^{k}q^{-k}$. If we instead write $\epsilon=q^{-(r-1)}$ for some $r$, then this becomes $q^{-rk}$ for $t=2^{d^{1+o(1)}}r^{1+o(1)}$. Apply Theorem 2.5 to find a $t$-rank-regular decomposition of $\mathbf{P}$ of size
@@
Put $V=\operatorname{Sp}\mathbf{Y}$ and $k=\dim V$ ($\leq s$). Consider the following two subspaces of $V$:
$U=\operatorname{Sp}\{X\in V\,|\,\operatorname{rk}(X)<tk\}$
and
$U_{\epsilon}=\operatorname{Sp}\{X\in V\,|\,\operatorname{bias}(X)\geq\epsilon q^{-k}\}.$
@@
    $U_{\epsilon}\subseteq U$. Indeed, for $X\in U_{\epsilon}$, write $X=X^{\uparrow}+X^{\prime}$ (so $\deg(X^{\uparrow})=\deg(X)$ and $\deg(X^{\prime})<\deg(X)$), and observe that $\operatorname{rk}(X^{\uparrow})=\operatorname{rk}(X)\leq tk$ by Theorem 2.10 and our choice of $t$. Thus, $X=X^{\uparrow}+X^{\prime}\in U+U=U$, using Lemma 2.6.
@@ 3 Applications
In this section we show that one can deduce strong bounds on the algebraic structure of polynomials, and polynomial maps, based solely on their image. The proofs use the weak regularity lemma.
We begin with the following quantification of algebraic structure, which constraints the degrees of the polynomials in the decomposition.
@@
This definition was used by Karam [15], and is closely related to an earlier definition of Green and Tao [10], $\operatorname{rank}_{t}(P)$, which expresses $P$ as any function of $\operatorname{rank}_{t}(P)$ polynomials of degree at most $t$ (over a finite field). Note that $\operatorname{rank}_{t}(P)\leq t\operatorname{rk}_{t}(P)$ (as each $t$-reducible can be expressed in terms of at most $t$ polynomials); thus, proving an upper bound on $\operatorname{rk}_{t}(P)$, as we will do below, would imply a similar upper bound on $\operatorname{rank}_{t}(P)$.
@@
If $\operatorname{rk}_{t}(P)$ is small, for $t$ bounded away from $\deg(P)$, then $P$ is “close” to having a considerably smaller degree than its actual degree (unlike most polynomials of a given degree). More precisely, by a standard Fourier-analytic argument (see e.g. [14]), $P\in\operatorname{Poly}(\mathbb{F}_{q})$ has correlation $|\operatorname{bias}(P-Q)|\geq q^{-\operatorname{rank}_{t}(P)}$ ($\geq q^{-t\operatorname{rk}_{t}(P)}$) with some polynomial $Q$ of degree at most tt. [^16] In the context of the Gowers inverse theorem, this is sometimes denoted $\left\lVert\omega^{P}\right\rVert_{u^{t+1}}\geq q^{-\operatorname{rank}_{t}(P)}$.
@@
Theorem 3.3 can be viewed as a result about systems of polynomials equations. Note that for any nonzero system of degree-1 polynomials $\mathbf{P}=(P_{1},\ldots,P_{m})$, there is a line [^17] In fact, an affine subspace of dimension $\dim\operatorname{Sp}(\mathbf{P})$. $\ell\in\mathbb{F}^{m}$ such that each system of equations $\mathbf{P}(\mathbf{x})=y$ for $y\in\ell$ has a solution (indeed, take any nonzero component $P_{i}$ and note that it must be surjective). Theorem 3.3 implies that for a system of (non-homogeneous) polynomials $\mathbf{P}$ of any degrees, if the above property doesn’t hold, then $\mathbf{P}$ must have a strong algebraic structure (namely, small $\operatorname{rk}_{d/2}(\mathbf{P})$).
@@
In fact, just the special case of the definition of $\operatorname{udeg}(P)$ where $(n,m,\operatorname{udeg}(P))=(1,1,1)$ is already very interesting. Indeed, note that for any finite field $\mathbb{F}$, a univariate polynomial $P\colon\mathbb{F}\to\mathbb{F}$ satisfies $\operatorname{udeg}(P)=1$ if and only if $P$ is a permutation of $\mathbb{F}$; such polynomials are naturally called permutation polynomials, and have attracted much attention (see e.g. [13] and the references within).
@@
We highlight two extreme special cases. One is where $\operatorname{udeg}(\mathbf{P})$ is very large, namely $\operatorname{udeg}(\mathbf{P})>\deg(\mathbf{P})/2$, in which case we directly obtain a small (depth-$3$) arithmetic formula for $\mathbf{P}$: each component $P_{i}$ is expressed as a linear combination of (the same) few products of degree-$1$ polynomials.

### v2

@@ 1.3 The proof
Towards the proof of the weak regularity lemma, we obtain results on linear spaces of polynomials and forms (homogeneous polynomials) that may be of independent interest. In particular, we analyze subspaces of high rank polynomials (where the rank is allowed to grow with the number of spanning polynomials, i.e., the dimension of the space). More particularly, we analyze high-rank polynomial “pencils”. By a polynomial pencil we mean a set of the form $V\setminus U$ where $V$ is a linear space of polynomials and $U$ is a strict subspace. For comparison, the classical regularity lemmas for polynomials work only with the “trivial” pencil $V\setminus\{0\}$; or more concretely, all nonzero linear combinations of the parts $\mathbf{X}$ are required to have high rank. Obtaining a high-rank pencil, rather than a high-rank trivial pencil, turns out to be achievable with far better bounds. We call this result a “rank-regularity lemma”, and note that its strong bounds eventually translate to strong bounds for our weak regularity lemma. Using it, the rest of our proof has, very roughly, the following logical structure (see Section 2 for all definitions):
$\mathbf{X}\text{ $t$-rank-regular}$ $\Leftrightarrow\operatorname{rk}(X_{1}+\operatorname{Sp}\mathbf{X}^{\prime})>t|\mathbf{X}|$
$\Rightarrow|\operatorname{bias}(X_{1}+\operatorname{Sp}\mathbf{X}^{\prime})|<\epsilon q^{-|\mathbf{X}|}$
@@
The proof moreover guarantees the additional required conditions, e.g., that $X_{1}$ has maximal degree. To translate from high rank to low bias as above, our proof relies on recent results on the relationship between structure and randomness for polynomials (see e.g. [Jan20, Mil19, CM21, MZ24]). Incidentally, these results are proved by going through the corresponding multilinear forms—which requires the characteristic of the field to be larger than the degree of the parts—by relating analytic rank and partition rank (defined by Gowers and Wolf [GW11] and by Naslund [Nas20], respectively).
@@ 1.4 Organization
In Section 2 we state and prove our weak regularity lemma. The proof has several components. In Section 2.1 we prove a rank-regularity lemma, which finds a small decompositions where every linear combination outside of a strict subspace has high rank. In Section 2.2 we show that a similar condition in terms of bias (rather than rank) suffices to get weak regularity, using Fourier-analytic techniques and known structure-vs-randomness results. Finally, in Section 2.3, we combine the aforementioned pieces to prove the weak regularity lemma. In Section 3 we obtain the applications of the weak regularity lemma outlined above.
@@ 1.5 Preliminaries
We will frequently use boldface letters to emphasize that an object is a tuple; so for example, we will use $\mathbf{P}$, instead of $P$, when dealing with a tuple of polynomials (or a polynomial map). We will use the convention that for a tuple $\mathbf{X}=(X_{1},\ldots,X_{k})$ and a set $A$, the notation $\mathbf{X}\subseteq A$ is a shorthand for $X_{i}\in A$ for all $i$. For $\mathbf{P}\subseteq\operatorname{Poly}$ we denote $\deg(\mathbf{P})=\max_{i}P_{i}$. More generally, for a set $V\subseteq\operatorname{Poly}$ we denote $\deg(V)=\max_{P\in V}\deg(P)$.
@@ 2 Weak Regularity
Definition 2.1 (weak $\epsilon$-regularity).
Let $\mathbb{F}=\mathbb{F}_{q}$. We say that $\mathbf{X}\in\operatorname{Form}^{k}(\mathbb{F})$ is weak $\epsilon$-regular if $\deg(X_{1})=\deg(\mathbf{X})$, and for every $y\in\mathbb{F}^{k}$,
@@
Let us also note that the probabilistic condition in Definition 2.1 means that $\Pr[\mathbf{X}=\mathbf{y}]/q^{-k}=\alpha\pm\epsilon$ for every $\mathbf{y}\in\mathbb{F}^{k}$ with $\Pr[\mathbf{X}^{\prime}=\mathbf{y}^{\prime}]/q^{-(k-1)}=\alpha$ (recall $\mathbf{X}^{\prime}$ is (k1)(k-1)-tuple). Intuitively, this means $\mathbf{X}$ is distributed like $\mathbf{X}^{\prime}$ up to an additive error of $\epsilon$.
@@ 2.1 Rank-regularity
An important definition in the study of polynomials is that of rank, sometimes also called strength [AH20] or Schmidt rank [Sch85]. Recall that a polynomial is reducible if it can be factorized as the product of two non-constant polynomials (all polynomials over the same field).
@@
Every $\mathbf{P}\in\operatorname{Poly}^{m}_{d}$ has a $t$-rank-regular decomposition of size at most $((2t+1)dm)^{2^{d}}$.
@@ 2.1.1 Linear spaces of polynomials
$V^{\uparrow}=\{X\in V|\deg(X)=\deg(V)\},$
where we recall the notation $\deg(V)=\max_{X\in V}\deg(X)$.
We will need the following closure property of the subspaces $U_{R}(V)$.
@@ 2.1.2 Proof of the rank-regularity lemma
Let $\mathbf{X^{0}}\in\operatorname{Form}^{r_{0}}$ be obtained from $\mathbf{P}\in\operatorname{Poly}^{m}$ by replacing each component of $\mathbf{P}$ with its homogeneous parts of degree $1,2,\ldots,d$. Note that $\mathbf{P}\subseteq\mathbb{F}[\mathbf{X^{0}}]$ and $r_{0}\leq dm$.
@@
For each $i$ starting with $i=0$, given $\mathbf{X^{i}}$ we proceed as follows:
  • Consider a minimal subspace $W$ of $\operatorname{Sp}(\mathbf{X^{i}})$, with a basis of forms $\mathbf{Y}\in\operatorname{Form}^{r^{\prime}}$, satisfying $\mathbf{P}\subseteq\mathbb{F}[W]$. Then $\mathbf{P}\subseteq\mathbb{F}[\mathbf{Y}]$ is a minimal decomposition, and moreover, $\deg(\mathbf{Y})\leq\deg(W)\leq\deg(\mathbf{X^{i}})$, and $r^{\prime}=\dim W\leq\dim\operatorname{Sp}(\mathbf{X^{i}})\leq r_{i}$.
  • Now, if $\mathbf{P}\in\mathbb{F}[\mathbf{Y}]$ is $t$-rank-regular then set $\mathbf{X^{i}}=\mathbf{Y}$ and $s=i$, finishing the construction (so $\mathbf{X^{s}}\in\operatorname{Form}^{r_{s}}$). Otherwise, $\mathbf{Y}$ is not $t$-rank-regular. We will use $\mathbf{Y}$ to construct $\mathbf{X^{i+1}}$ with $\deg(\mathbf{X^{i+1}})<\deg(\mathbf{X^{i}})$. Recall that in order to apply Lemma 2.9, we need a tuple of forms satisfying two conditions: they are all of the same degree, and the tuple is not $t$-rank-regular.
    Put $\ell=\deg(\mathbf{Y})$, and let $\mathbf{Y^{*}}$ be obtained from $\mathbf{Y}$ by replacing every component polynomial of degree smaller than $\ell$ by an arbitrary component polynomial of degree $\ell$; that is, each component of $\mathbf{Y}$ that lies outside of $\mathbf{Y}^{\uparrow}$ is replaced in $\mathbf{Y^{*}}$ by a polynomial from $\mathbf{Y}^{\uparrow}$. [^13] That is, $\mathbf{Y^{*}}$ is obtained from $\mathbf{Y}^{\uparrow}$ by arbitrarily duplicating components to achieve size $r^{\prime}$. Note that $\mathbf{Y^{*}}\in\operatorname{Form}^{r^{\prime}}_{\ell}$ (whereas $\mathbf{Y}\in\operatorname{Form}^{r^{\prime}}$). We claim that $\mathbf{Y^{*}}$ is not $t$-rank-regular, so that we would be able to apply Lemma 2.9 on it. We have that $\mathbf{Y^{*}}$ and $\mathbf{Y}^{\uparrow}$ are equal as sets: every component of $\mathbf{Y^{*}}$ lies in $\mathbf{Y}^{\uparrow}$, and every component of $\mathbf{Y}^{\uparrow}$ lies in $\mathbf{Y^{*}}$ (though $\mathbf{Y^{*}}$ has the same size as $\mathbf{Y}$). Put $V=\operatorname{Sp}\mathbf{Y}$. Since $\mathbf{Y}$ is not $t$-rank-regular, we have $U_{R}(V)=V$ for $R=tr^{\prime}$. By Corollary 2.8, $U_{R}(V^{\uparrow})=V^{\uparrow}$. This implies that $\mathbf{Y^{*}}$ is not $t$-rank-regular, as claimed, using that $V^{\uparrow}=(\operatorname{Sp}\mathbf{Y})^{\uparrow}=\operatorname{Sp}(\mathbf{Y}^{\uparrow})=\operatorname{Sp}(\mathbf{Y^{*}})$ and $|\mathbf{Y^{*}}|=|\mathbf{Y}|$ (that is, $R=t|\mathbf{Y^{*}}|$).
    Apply Lemma 2.9 on $\mathbf{Y^{*}}$. Merge the resulting tuple of forms with the forms in $\mathbf{Y}$ of degree smaller than $\ell$ to obtain $\mathbf{X^{i+1}}\in\operatorname{Form}^{r_{i+1}}$ with $\mathbf{Y}\subseteq\mathbb{F}[\mathbf{X^{i+1}}]$, $\deg(\mathbf{X^{i+1}})<\deg(\mathbf{X^{i}})$, and $r_{i+1}\leq 2tr_{i}^{2}+r_{i}\leq(2t+1)r_{i}^{2}$.
To prove that this construction gives the claim above, we proceed by induction on $i\geq 0$. The induction base $i=0$ follows trivially from the construction of $\mathbf{X^{0}}$, as $\mathbf{P}\subseteq\mathbb{F}[\mathbf{X^{0}}]$ with $\deg(\mathbf{X^{0}})\leq d$, and $r_{0}\leq dm=(2t)^{0}(dm)^{1}$. For the induction step we have, by the construction of $\mathbf{X^{i+1}}$ and the induction hypothesis, that $\mathbf{P}\subseteq\mathbb{F}[\mathbf{Y}]\subseteq\mathbb{F}[\mathbf{X^{i+1}}]$, that $\deg(\mathbf{X^{i+1}})\leq\deg(\mathbf{X^{i}})-1\leq d-(i+1)$, and that $r_{i+1}\leq(2t+1)r_{i}^{2}\leq(2t+1)((2t+1)^{2^{i}-1}(dm)^{2^{i}})^{2}=(2t+1)^{2^{i+1}-1}(dm)^{2^{i+1}}$, as needed. This completes the proof. ∎
@@ 2.2 Bias and weak regularity
We note that $|\operatorname{bias}(P)|$ is closely related to the statistical distance between the distribution of $P$ and the uniform distribution on $\mathbb{F}$ (see e.g. [BV10], Claim 33).
@@
Theorem 2.10 ([MZ24], Corollary 8.5141414See also Remark 8.3 for the bound on $c_{d}$.).
@@
We also remark that one can replace the use of Theorem 2.10 in our proof by earlier bounds of Milićević [Mil19] or Janzer [Jan20], and still obtain reasonable, albeit somewhat worse bounds in Theorem 2.2.
@@ 2.3 Putting everything together
Let $t$ be such that the bound in Theorem 2.10 satisfies $q^{-c_{d}tk/\mathcal{L}_{\mathbb{F}_{q}}(tk)}<\epsilon q^{-k}$ for every $k\leq((2t+1)dm)^{2^{d}}$ (recall $\epsilon=q^{-r})$. We observe that $t=2^{d^{1+o(1)}}(r+1)^{1+o(1)}\log(m)$ suffices. [^17] In fact, this choice of $t$ suffices even if the right hand side is $\epsilon^{k}q^{-k}$. If we instead write $\epsilon=q^{-(r-1)}$ for some $r$, then this becomes $q^{-rk}$ for a similar $t$. Apply Theorem 2.5 to find a $t$-rank-regular decomposition of $\mathbf{P}$ of size
@@
Put $V=\operatorname{Sp}\mathbf{Y}$ and $k=\dim V$ ($\leq s$). Consider the subspace of $V$,
$U=\operatorname{Sp}\{X\in V\,|\,\operatorname{rk}(X)<tk\},$
and subset of $V$,
$U_{\epsilon}=\{X\in V\,|\,\operatorname{bias}(X)\geq\epsilon q^{-k}\}.$
@@
    $U_{\epsilon}\subseteq U$. Indeed, for $X\in U_{\epsilon}$ we have $\operatorname{rk}(X)<tk$ by Theorem 2.10 and our choice of $t$.
@@ 3 Applications
In this section we obtained several results along the theme that one can deduce strong bounds on the algebraic structure of polynomials, and polynomial maps, based solely on their image. The proofs use the weak regularity lemma.
We begin with the following standard quantification of algebraic structure, which constraints the degrees of the polynomials in the decomposition.
@@
This definition was used by Karam [Kar23], and is closely related to an earlier definition of Green and Tao [GT09], $\operatorname{rank}_{t}(P)$, which expresses $P$ over a finite field as any function of $\operatorname{rank}_{t}(P)$ polynomials of degree at most $t$. (Note that $\operatorname{rank}_{t}(P)\leq t\operatorname{rk}_{t}(P)$, as each $t$-reducible can be expressed in terms of at most $t$ polynomials; thus, proving an upper bound on $\operatorname{rk}_{t}(P)$, as we will do below, would imply a similar upper bound on $\operatorname{rank}_{t}(P)$.)
@@
If $\operatorname{rk}_{t}(P)$ is small, for $t$ bounded away from $\deg(P)$, then $P$ is “correlated” with a polynomial having a considerably smaller degree (unlike most polynomials of a given degree). More precisely, by a standard Fourier-analytic argument (see e.g. [Jan20]), $P\in\operatorname{Poly}(\mathbb{F}_{q})$ has correlation $|\operatorname{bias}(P-Q)|\geq q^{-\operatorname{rank}_{t}(P)}$ ($\geq q^{-t\operatorname{rk}_{t}(P)}$) with some polynomial $Q$ of degree at most tt. [^18] In the context of the Gowers inverse theorem, this is sometimes denoted $\left\lVert\omega^{P}\right\rVert_{u^{t+1}}\geq q^{-\operatorname{rank}_{t}(P)}$.
@@
Theorem 3.3 can be viewed as a result about systems of polynomials equations. Note that for any $\mathbf{P}$ consisting of homogeneous polynomials of degree $d$ over $\mathbb{F}_{q}$ with $\gcd(d,q-1)=1$, if the system of equations $\mathbf{P}(x)=y_{0}$ has a solution for some $y_{0}$, then so does $\mathbf{P}(x)=y$ for each $y$ on the line through $y_{0}$ and 0 (indeed, if $\mathbf{P}(x)=y_{0}$ then $\mathbf{P}(cx)=c^{d}\mathbf{P}(x)=ty_{0}$ for $t=c^{d}$). Thus, either $\mathbf{P}(x)=y$ has a solution for each $y$ on some line, or else $\mathbf{P}\equiv 0$. Theorem 3.3 implies that for arbitrary polynomials $\mathbf{P}$ of any degrees, either the system $\mathbf{P}(x)=y$ has a solution for each $y$ on some line, or else $\mathbf{P}$ has a strong algebraic structure (namely, small $\operatorname{rk}_{d/2}(\mathbf{P})$).
@@
The special case of Definition 3.5 where $(n,m,\operatorname{udeg}(P))=(1,1,1)$ is already very interesting. Indeed, note that for any finite field $\mathbb{F}$, a univariate polynomial $P\colon\mathbb{F}\to\mathbb{F}$ satisfies $\operatorname{udeg}(P)=1$ if and only if $P$ is a permutation of $\mathbb{F}$; such polynomials are naturally called permutation polynomials, and have attracted much attention (see e.g. [Hou15] and the references within).
@@
We highlight two extreme special cases. One is where $\operatorname{udeg}(\mathbf{P})$ is very large, namely $\operatorname{udeg}(\mathbf{P})>\deg(\mathbf{P})/2$, in which case we directly obtain a small depth-$3$ (homogeneous) arithmetic formula for $\mathbf{P}$: each component $P_{i}$ is expressed as a linear combination of (the same) few products of degree-$1$ polynomials.